OpenStudy (anonymous):

A student pushes a 40-N block across the floor for a distance of 10 m. How much work was applied to move the block? A. 4 J B. 40 J C. 400 J D. 4000 J

OpenStudy (timaashorty):

I'm not 100% sure about this one but I would go with C.

OpenStudy (anonymous):

I think it's also C because work=force x distance.

OpenStudy (theeric):

Right! Because the force and displacement are in the same exact direction, you can use that formula.

OpenStudy (anonymous):

thanks everyone! You all are very helpful!

OpenStudy (theeric):

:) You're welcome! More generally, \(\large W=F\dot\ d\dot\ \cos(\theta)\) where \(\theta\) is the angle between force and displacement. You can see that, when they're in the same direction, \(\theta=0\) \(\implies W=F\dot\ d\dot\ \cos(0)=F\dot\ d\dot\ 1=F\dot\ d\)